nonbinarycode:

In Circle O, AB = 12 centimeters and mAOB = 60°. Sin 60° is of 360° then the length of is the circumference of the circle. Use this fact to find the exact length of . 4(pi) cm 24(pi) cm 4 cm 8(pi) cm

nonbinarycode:

mhanifa:

Do you know the circumference formula?

mhanifa:

We know the circle is 360 deg. We have an arc AB of 60 deg which is 12 cm. Circumference is: \[C = 2\pi*r\]

mhanifa:

And we have: \[AB = (2\pi*r)*60/360 = 12 cm\]

mhanifa:

And what is the exact question and your plan?